A scalene trinagle with the height of 4cm and the base of 6cm?
Answer by jim_thompson5910(35256)

If you want to find the area of this triangle, then the area is

A = (bh)/2

A = (6*4)/2

A = 24/2

A = 12


So the area of this triangle is 12 square centimeters.
--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------